OpenStudy (anonymous):

PLEASE HELP!!!! What is the wavelength of an electromagnetic wave that travels at 3x10^8 m/s and has a frequency of 60 MHz?

OpenStudy (abb0t):

\(\sf \color{}{300 \times 10^6 \frac{meters}{sec}\times \frac{cycles}{60 \times 10^6~sec}}\) which is meters per cycle, and that is ur wavelength.